Y - 0.72 Y = 100 + 200 + 180

0.28Y = 480. how u get 0.28?

Grade:12th pass

2 Answers

Harshit Porwal
7 years ago
Hello, I will help you.Actually there is no value given of `y` so we will take `y` as 1y and 0.72y as 0.72y. So,1y-0.72y=100+200+1800.28y=480y=1714.29Hope this will help you!Harshit Porwal
Harshit Porwal
7 years ago
Sorry i don`t know i have to use commands!
There is no value given of `y` so we will take `y` as 1y
1y-0.72y=100+200+180
0.28y=480
y=1714.29